OpenStudy (anonymous):

The sum of the angle measures of a polygon with s sides is 2520. Find s. Please explain this to me!

OpenStudy (anonymous):

here im assuming that you meant the sum of interior angles in a polygon add upto 2520. if thats the case you need to know the general rule: (n-2)*180 = sum of interior angles where n = number of sides. so: (n-2)*180 = 2520---> since this is what it adds upto. (n-2) = 2520/180 n = 14 + 2 so n...or in your case 's' = 16